OpenStudy (anonymous):

write .25x^2-.5x-5 in vertex form

OpenStudy (anonymous):

do you know what the vertex is?

OpenStudy (anonymous):

nope another question is to find vertex

OpenStudy (anonymous):

Vertex form: y=a(x-h)2+k (h,k) is the vertex.

OpenStudy (anonymous):

so how do i rearrange it? thank you

OpenStudy (anonymous):

easiest to find the vertex first then write in vertex form do you know how to find the vertex of \[y=ax^2+bx+c\]?

OpenStudy (anonymous):

im not sure no

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

You already have the equation in that form. .25x^x - .5x -5

OpenStudy (anonymous):

no the vertex form i need to change it from that form to vertex do i have to complete square?

OpenStudy (anonymous):

You have to find (h,k) which is the vertex. If you want, you can just use a graphing calculator

OpenStudy (anonymous):

So the vertex is about (0, -5)

OpenStudy (anonymous):

Plug that into the equation

OpenStudy (anonymous):

oh no !!

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{2a}\) which in your case is \[-\frac{-.5}{2\times .25}=1\]

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

Oh, thanks for correcting me

OpenStudy (anonymous):

to find the second coordinate, replace \(x\) by \(1\)

OpenStudy (anonymous):

THANK YOU GUYS

OpenStudy (anonymous):

you are not done yet though

OpenStudy (anonymous):

Okay, after you get the coordinates, you plug them into y=a(x-h)2+k

OpenStudy (anonymous):

h is x and k is y (h, k) = (x, y)

OpenStudy (anonymous):

\[y=.25x^2-.5x-5\] \[y=.25-.5-5=-5.25\] so vertex form is \[y=.25(x-1)^2-5.25\]

OpenStudy (anonymous):

yay great job thank you thank you thank you!!!! also x intercepts are when you set the equation to 0 and solve?

OpenStudy (anonymous):

or when you fill in 0 for x?

OpenStudy (anonymous):

It's when you make y=0

OpenStudy (anonymous):

To find the y intercept, you make x=0

OpenStudy (anonymous):

i appreciate you all immensely thank you!!!!

OpenStudy (anonymous):

Good job with doing the work! :)